OpenStudy (anonymous):

Solve. 2x^2 + 16x + 50 = 0 a. -4 +- i b.-4 +- 3i c.-2+-4i d. -3+-4i

hero (hero):

Complete the square

OpenStudy (anonymous):

idkk how to do thatt........

hero (hero):

2x^2+16x+50 = 0 2(x^2+8x+25) = 0 x^2+8x+25 = 0 x^2+8x = -25 x^2+8x+16 = -25 + 16 (x+4)^2 = -9 x + 4 = sqrt(-9) x = sqrt(-9) -4 x = plus or minus 3i -4 Option b